OpenStudy (anonymous):

Can someone help me with this question? 17. what is the measure of angle B in the triangle below? a. 21.0 degrees b. 22.6 degrees c. 42.7 degrees d. 67.4 degrees

Mehek (mehek14):

triangle image?

OpenStudy (anonymous):

Could someone help explain it also, because I honestly don't know how to do it.

OpenStudy (campbell_st):

well you need use a trig ratio.... it doesn't matter which ratio you select as they will all give the same answer... any choices..?

OpenStudy (anonymous):

I don't know what you mean.

OpenStudy (campbell_st):

have you studied trigonometry...?

OpenStudy (alexandervonhumboldt2):

\(\huge\color{lime}{{\rm Welcome}\rm~to~Open~Study!!!!}\) @BrinleeMermaid \(\Huge \color{blue}{\star^{ \star^{\star:)}}}\Huge \color{red}{\star^{ \star^{\star:)}}}\) \(\Huge \color{green}{\star^{ \star^{\star:)}}}\Huge \color{yellow}{\star^{ \star^{\star:)}}}\)

OpenStudy (jdoe0001):

do you know what the "opposite side" for angle B is in that triangle?

OpenStudy (jdoe0001):

anyhow.... all you'd need is to recall your SOH CAH TOA \(\bf sin(\theta)=\cfrac{opposite}{hypotenuse} \qquad \qquad % cosine cos(\theta)=\cfrac{adjacent}{hypotenuse} \\ \quad \\ % tangent tan(\theta)=\cfrac{opposite}{adjacent}\)

OpenStudy (anonymous):

I know that, but it isn't the side that we are trying to find, it's the angle.

OpenStudy (jdoe0001):

right so.. ahemm say... do you know what the "opposite side" of angle B is?

OpenStudy (jdoe0001):

keep in mind, an opposite side, is the one "facing off" the angle so if you put your "eye" on that angle, the side you'd be facing will be the "opposite side" so.... which is it from angle B then?

OpenStudy (anonymous):

the left side

OpenStudy (anonymous):

hello?

OpenStudy (jdoe0001):

in this case, yes, or side 5 thus  so that means \(\bf sin(\theta)=\cfrac{opposite}{hypotenuse}\implies sin^{-1}[sin(\theta)]=sin^{-1}\left( \cfrac{opposite}{hypotenuse} \right) \\ \quad \\ \theta=sin^{-1}\left( \cfrac{opposite}{hypotenuse} \right) \\ \quad \\ % cosine cos(\theta)=\cfrac{adjacent}{hypotenuse}\implies \implies cos^{-1}[cos(\theta)]=cos^{-1}\left( \cfrac{adjacent}{hypotenuse} \right) \\ \quad \\ \theta=cos^{-1}\left( \cfrac{adjacent}{hypotenuse} \right) \\ \quad \\ % tangent tan(\theta)=\cfrac{opposite}{adjacent}\implies \implies tan^{-1}[tan(\theta)]=tan^{-1}\left( \cfrac{opposite}{adjacent} \right) \\ \quad \\ \theta=tan^{-1}\left( \cfrac{opposite}{adjacent} \right)\)

OpenStudy (anonymous):

to find the degree of the angle?

OpenStudy (jdoe0001):

yes, you could use any of them really because you're given all sides so use either, sine or cosine or tangent

OpenStudy (anonymous):

that still doesn't make sense to me. there should be a certain formula to find the degree of the angle. you confused me by putting all of that mumbo jumbo up there.

OpenStudy (jdoe0001):

well... have you covered \(\bf sin(\theta)=\cfrac{opposite}{hypotenuse} \qquad \qquad % cosine cos(\theta)=\cfrac{adjacent}{hypotenuse} \\ \quad \\ % tangent tan(\theta)=\cfrac{opposite}{adjacent}\) yet?

OpenStudy (anonymous):

yes, I already said that. could you explain to me how to find the degree of any triangle using those??

OpenStudy (jdoe0001):

welll sure, let us use say the sine then \(\bf sin(\theta)=\cfrac{opposite}{hypotenuse}\implies sin^{-1}[sin(\theta)]=sin^{-1}\left( \cfrac{opposite}{hypotenuse} \right) \\ \quad \\ \theta=sin^{-1}\left( \cfrac{opposite}{hypotenuse} \right)\) the \(\Large sin^{-1}\) should be in your calculator, so just plug in the values for opposite and hypotenuse make sure your calculator is in Degree mode

OpenStudy (jdoe0001):

opposite is 5 hypotenuse 15 thus \(\bf sin(B)=\cfrac{5}{13}\implies sin^{-1}[sin(B)]=sin^{-1}\left( \cfrac{5}{13} \right) \\ \quad \\ \measuredangle B=sin^{-1}\left( \cfrac{5}{13} \right)\)

OpenStudy (jdoe0001):

hmm opposite is 5 hypotenuse 13 anyway

OpenStudy (anonymous):

okay I got it now. you weren't explaining it very well at first.

OpenStudy (jdoe0001):

heheh, ok